OpenStudy (anonymous):

how to reduce 316 over 126 i know its improper but i don't kno what to next

OpenStudy (turingtest):

316/126 if in doubt just divide top and bottom by 2 at a time because both are even 316/126=158/63 then try to factor top and bottom 158/63=(2x79)/(3^2x7) so it looks like we can't reduce it any further.

OpenStudy (anonymous):

Well if you do this 316/126 divide it bye 2 then you can keep going down if you need to so the answer would be....158/63 then you divide those to from each other and you get.....2 1/2 and thats your answer....